Yan Yan Corp. has a $4,000 par value bond outstanding with a coupon rate of 4.9 percent paid semiannually and 11 years to maturity. The yield to maturity on this bond is 4.2 percent.

What is the price of the bond? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

HI,

Here par value of bond F = $4000

coupon rate = 4.9% paid semiannually

so per period coupon= 4000*4.9%/2 = $98

time t = 11 years =22 semi annual

YTM = 4.2% = 2.1% semi annual

Price of bond = C*(1-(1/(1+YTM))^t/YTM +4000*(1/(1+YTM)^t)

Price of bond = 98*(1-(1/1+2.1%)^22/0.021 + 4000*(1/1+2.1%)^22

Price of bond = $4,244.64

Thanks